You have deposited $10,800 into an account that will earn an interest rate of 15% compounded semiannually. How much will youHow do I calculate this on a financial calculator?
Answer #1

We use the formula:
A=P(1+r/2)^2n
where
A=future value
P=present value
r=rate of interest
n=time period.

A=10800*(1+0.15/2)^(2*4)

=10800*1.783477826

=$19261.56(Approx).